OpenStudy (anonymous):

With the numbers 8, 15, and 22, adding, subtracting, and dividing only, how do you make the solution end up being 9?

OpenStudy (anonymous):

@bubblysmiles24

OpenStudy (anonymous):

8+15 =23 23-22=1 1+15=16 16/8=2 2+15=17 17-8=9

OpenStudy (rational):

\[8+8+8-15 = 9\]

OpenStudy (rational):

does that work too ? or we need to use division also

OpenStudy (anonymous):

Dang.... That's a lot easier...

OpenStudy (rational):

Haha i like ur solution though, next step follows from present step

OpenStudy (dan815):

8 , 15 :)